OpenStudy (anonymous):

3 minus n over 2 = 3 over 4 to do this i cross multiply right or wrong

OpenStudy (anonymous):

And then you have to substract and the divide

OpenStudy (anonymous):

i did that and i get 3 over 2 is that right

OpenStudy (anonymous):

3-n = 2(3/4) 3-n=6/4 4(3-n)=6 12-4n=6 -4n=6-12 -4n=-6 n=-6/-4 n=3/2

OpenStudy (anonymous):

thank you